OpenStudy (anonymous):

A tugboat goes 180 mils upstream in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current. The speed of the tugboat is ______ mph, and the speed of the current is _______mph. (Simplify your answers).

OpenStudy (anonymous):

We know: \[ \begin{align} \frac{180}{10}&=18\text{ mph}\\ \frac{180}{5}&=36\text{ mph} \end{align} \]Since the current is constant, then, we know it goes: \[ \frac{36-18}{2}=9\text{ mph} \]Therefore the speed of the tugboat is: \[ \frac{18+36}{2}=18+9=36-9=27\text{ mph} \]Without a current. And, by this, we are done.

OpenStudy (anonymous):

@LolWolf Thank you for not just giving me the answer but showing the work. It helps me so that I am able to solve the problem myself next time. Thank you, thank you thank you. One more question. How do I know when to divide?

OpenStudy (anonymous):

Sure thing. And, we know that the speed of the boat is the average of the speeds going upstream and downstream. (i.e. if the boat goes at a speed \(v\) and the river goes at a speed \(r\), then, it goes up-stream at a speed \(u=v-r\) and downstream at a speed \(d=v+r\), which means, that, solving for \(v\): \[ \frac{d+u}{2}=\frac{v+r+v-r}{2}=\frac{v+v+r-r}{2}=\frac{2v}{2}=v \]And the river's speed is: \[ \frac{d-u}{2}=\frac{v+r-(v-r)}{2}=\frac{v+r-v+r}{2}=\frac{v-v+r+r}{2}=\frac{2r}{2}=r \]Since we are given \(d\) and \(u\) we use the formulas to show this.

OpenStudy (anonymous):

Thank you!!!